A 123.1 g sample of water at 25.5°C is heated to steam at 126.5°C. How much heat was absorbed?
1
Expert's answer
2019-06-10T09:45:01-0400

The total heat equals to

"Q=0.1231(74.5\cdot4.2\cdot10^{3} +2.26\cdot10^{6}+ 4\cdot8.31\frac{10^3}{18}26.5)"

i.e.

"Q = 0.32 (MJ)."